OpenStudy (anonymous):

One number is five more than three times another. If their difference is 35, what is the larger number?

OpenStudy (anonymous):

50/?

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

let two no. are "X" and "Y" as "X">"Y" ;then "X-3Y=5"first equation "X-Y=35" second equation by solving 2Y=30 then "Y"=15 ;then "X"=15+35=50